OpenStudy (anonymous):

In the first set of problems. Units and Dimensional Analysis Challenge Problem Solutions - Problem 1: How did they find (b/g)^1/2 ?

OpenStudy (anonymous):

please give the link of question.

OpenStudy (anonymous):

-3V+W+X+2(Y+Z)=0 Then, on solving other two equation we get V=0 x=-1/2 on substituting it in first equation we get W+(-1/2)+2(Y+Z)=0 or, W+2(Y+Z)=1/2 ................................ (1) Now since from equation of continuity we have \[hA_1=vA_2t\] or \[t=\frac{h}{v} \frac{A_1}{A_2}\] from here we get power of A1 is 1 and power of A2 is -1. hence y=1 and z=-1 put it in equation (1) we get w=1/2.

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

you are most welcome.